# A Survey on Collaborating Small and Large Language Models for Performance, Cost-effectiveness, Cloud-edge Privacy, and Trustworthiness

### v1

@@ 4.2 Fine-tuning Stage
During fine-tuning, edge data can enhance both edge and cloud LLMs, though protecting sensitive information remains challenging. Depending on the enhanced target, privacy-preserving edge–cloud frameworks typically adopt three paradigms: (1) SLMs as Learners, gaining knowledge from LLMs through local fine-tuning or distillation; (2) LLMs as Learners, improving via proxy signals from tuned SLMs; and (3) Collaborative Learners, where both models co-evolve through federated or privacy-preserving knowledge exchange.
@@
(i) Raw Data: SLMs enhance reasoning by learning from cloud LLM outputs while keeping private data local. The LLM provides interpretable supervision (e.g., responses, reasoning traces) or synthetic samples for SLM fine-tuning. MiniLLM Gu et al. (2024) and LlamaDuo Park et al. (2025) distill reasoning traces, while DRAG Chen et al. (2025b) extends to retrieval-augmented reasoning. Privacy-aware frameworks such as HomeLLaMA Huang et al. (2025b) and Qin et al. (2024) use selective textual supervision to align SLMs with user preferences under strict privacy constraints.
(ii) Logit-based Learning: SLMs learn from teacher LLMs’ numerical outputs, using logits as soft targets to imitate their decision boundaries without accessing raw text. TinyLLM Tian et al. (2025) distills from multiple teachers for robustness, Mix Distillation Li et al. (2025d) combines LLM and SLM supervision, and ADPA Gao et al. (2025) introduces an advantage-based objective for better reasoning alignment, ensuring privacy and efficiency through abstracted signal exchange.
@@
(i) Logit-based Learning: Instead of directly fine-tuning LLMs, Ormazabal et al. (2023) steer them via SLMs trained on private data, exposing only logits for guidance, with later extensions using logit offsets Liu et al. (2024a); Wang et al. (2025d). Follow-up work addresses training–inference inconsistency He et al. (2024); Yao et al. (2025) and improves robustness through re-encoding SLM outputs Fang et al. (2024a), enabling adaptation across SLM–LLM pairs without shared vocabularies while preserving data privacy.
(ii) LoRA-based Learning: Recent work explores using LoRA signals for knowledge transfer. PHLoRA Vasani et al. (2025) avoids dual logit outputs by computing low-rank adaptations with strong performance gains. LoRASuite Li et al. (2025c) and LoRA-X Farhadzadeh et al. (2025) enable cross-version and cross-architecture migration, while Trans-LoRA Wang et al. (2024c) and Cross-LoRA Xia et al. (2025) generalize LoRA transfer across heterogeneous models for scalable, privacy-conscious LLM adaptation.
@@
Beyond one-way transfer, collaboration can be bidirectional: SLMs and LLMs co-adapt through shared representations, logits, or adapter weights while preserving privacy (Fig. 15). Edge SLMs train on local data and send abstracted updates (e.g., logits, LoRA signals) for cloud aggregation. CrossLM Deng et al. (2025) combines private data and LLM-generated samples for co-training, while LSRP Zhang et al. (2025b) enables mutual retrieval and response refinement. Federated frameworks Fan et al. (2025b); Gao et al. (2024); Peng et al. (2024); QI et al. (2024); Fang et al. (2024b) aggregate SLM updates to refine central LLMs, with FLoRA Wang et al. (2024d) enhancing personalization through improved edge aggregation.
@@ 5 Collaboration for Trustworthiness
LLMs show strong generative ability but face trustworthiness issues, including hallucinations, hazards, bias, privacy leaks, and jailbreak. Updating large models for safety is costly and inflexible. As a remedy, SLMs offer lightweight, adaptable external safety control. Recent collaborations adopt two forms: (1) Safety-guided decoding, where SLMs adjust LLM logits during generation, and (2) Guardian–generator, where SLMs filter inputs or outputs, ensuring safer LLM deployment.
@@ 5.1 Safety-Guided Decoding
Unsafe LLM outputs often emerge during decoding, making token-level safety control essential yet difficult to anticipate. Recent methods use safety-tuned SLMs to guide LLM decoding against jailbreaks, privacy leaks, and bias. As shown in Fig. 16, SLM and LLM logits are fused for safe, high-quality generation via (1) direct logit fusion or (2) offset fusion, where SLM logits steer LLM probabilities toward safer tokens.
@@
(1) Direct Logit Fusion The CP-$\Delta$ theory Vyas et al. (2023) provides an information-theoretic foundation for model fusion, demonstrating that combining the predictive distributions of two models, $q_{1}(y|x)$ and $q_{2}(y|x)$, with distinct copyright risks yields a fused distribution $p(y|x)\propto q_{1}(y|x)^{\alpha}q_{2}(y|x)^{1-\alpha}$ that minimizes weighted KL divergence and balances their strengths without retraining, where $\alpha$ controls each model’s contribution. The fused distribution lies closer to both sources than they are to each other, thus achieving a purification effect, retaining desirable properties while reducing copyrighted content. This operation is equivalent to linear weighting in logit space, $z_{p}=\alpha z_{1}+(1-\alpha)z_{2}$, providing the theoretical foundation for SLM–LLM collaborative logit fusion. Subsequent methods build on this principle: Purifying LLMs Li et al. (2024) fuses clean SLM and contaminated LLM logits to mitigate copyright, privacy, and backdoor risks, while MOD Shi et al. (2024) extends fusion to multiple objectives, integrating models optimized for fluency, faithfulness, and safety to enable controllable generation.
(2) Logit Offset Fusion Mitchell et al. (2024) introduced the concept of logit offset fusion by decomposing a fine-tuned model’s output into the base model’s logits and a behavioral delta. Through an up-scaling strategy, it integrates a large pre-trained model’s logits with the behavioral delta of a small fine-tuned model, achieving 73% of full large-model fine-tuning helpfulness without retraining. This demonstrates that safety-aligned small models’ behavioral deltas can enhance LLM safety when fused at the logit level. Building on this idea, Offset Unlearning Huang et al. (2025a) trains SLMs to generate targeted logit offsets that suppress sensitive or fabricated knowledge for privacy-preserving unlearning. Conversely, Weak-to-Strong Jailbreak Zhao et al. (2025) reveals that unsafe SLMs can steer safe LLMs toward harmful outputs, implying that the same mechanism can defensively employ safe SLMs to counter jailbreak attacks through safety-guided logit fusion.
@@ 5.2 SLM as Guardian, LLM as Generator
Figure 17: SLM Guardian, LLM Generator
Building on the SLM-as-Guardian paradigm, subsequent works further advance safety coverage and efficiency. The Llama Guard series Inan et al. (2023); Team (2024); Dubey et al. (2024); Fedorov et al. (2024); Chi et al. (2024); AI (2025a) evolves from basic LLaMA-based classifiers to multilingual, cross-modal, and domain-adapted guardrails with fine-grained risk taxonomies. Lightweight filters such as Prompt Guard AI (2024, 2025b), ShieldGemma Zeng et al. (2024, 2025), and WildGuard Han et al. (2024) improve efficiency and robustness. ThinkGuard Wen et al. (2025) enhances interpretability, MiniCheck Tang et al. (2024) supports factual verification, Cascade Guardrails Nagireddy et al. (2024) introduce hierarchical defenses, and LlamaFirewall Chennabasappa et al. (2025) enables multi-agent protection.
@@ 6 Challenges and Future Directions
More future directions are in Appendix A.2.
Open and Interoperable Ecosystem. Existing SLM–LLM collaborations are fragmented and task-specific. Future work should build open, cross-domain ecosystems with standardized interfaces and shared repositories to enable model reuse, composition, and joint benchmarking.
Comprehensive Benchmarking. Current evaluations emphasize open-ended performance but neglect structured and system-level metrics. Future benchmarks should assess structured reasoning, modular cooperation, and efficiency indicators such as throughput, latency, and cost–performance.
Unified Safety Collaboration. Present safety methods decouple decoding and moderation. Future research should unify them, allowing SLMs to act as safety-aware decoders and guardians for end-to-end, trustworthy generation and interaction.
@@ 7 Conclusions
Appendix A Appendix
A.1 Applications
Collaboration for performance In Guidance–Generation, SLMs aid LLMs in reasoning and classification (intent detection, NLI, sentiment, multi-hop QA). In Division–Fusion, they form pipelines for domains such as clinical analysis, factual verification, NL2SQL, and trajectory modeling, with SLMs handling subtasks and LLMs integrating results.
Collaboration for cost-effectiveness These show stage-specific application preferences. Inference-stage approaches emphasize real-time reasoning and conversational tasks where latency and compute efficiency are critical. Tuning-stage methods focus on intent classification and reasoning transfer, enabling smaller models to inherit LLM capabilities with minimal cost. Pre-training-stage frameworks target corpus construction and scaling efficiency on large datasets, supporting sustainable and scalable model development.
Collaboration for cloud-edge privacy Inference-stage solutions suit end-user scenarios with strict privacy needs, where edge SLMs prevent PII or sensitive context leakage in domains such as healthcare and customer service. Depending on privacy levels, SLMs can act as Sensitive Information Gatekeepers or All-Information Guardians. Meanwhile, proxy tuning enables domain adaptation of cloud LLMs without direct data exposure, offering broader applicability under edge resource limits and SLMs’ constrained reasoning capacity.
Collaboration for trustworthiness SLM–LLM collaboration for trustworthiness has been applied in safety-critical and compliance-sensitive contexts. Safe decoding methods mitigate hallucination, privacy, and toxicity risks in code generation, instruction following, and dialogue. Guardian frameworks support moderation, assistants, and fact verification by filtering unsafe inputs and ensuring output integrity. Together, these works show that SLMs function as lightweight, adaptive safety modules, embedded for fine-grained control or external for interactive oversight, enabling scalable, reliable AI deployment.
A.2 More Future Directions
Cost-Aware SLM–LLM Collaboration Current SLM–LLM collaboration paradigms mostly prioritize performance gains without a rigorous, comprehensive cost model, potentially leading to inefficient resource allocation where the operational expense negates the utility benefits. Future research can develop intelligent, adaptive orchestration frameworks that dynamically optimize the cost-performance trade-off. This includes investigating predictive cost-aware routing to direct queries to the most economically efficient model, designing non-myopic collaboration policies that consider long-term inference budgets, and creating multi-objective optimization techniques that balance accuracy, latency, and financial cost for sustainable and scalable SLM–LLM systems.
Theory for Privacy-Preserving Collaboration Current privacy-preserving SLM–LLM collaborations lack formal theoretical foundations, relying mostly on empirical evaluations without quantifiable privacy assurances. Future research should establish rigorous metrics and provable bounds for information leakage and privacy preservation, enabling verifiable guarantees of data protection and fostering trust in large-scale, privacy-sensitive SLM–LLM systems.
I. Performance Collaboration
[…]
Sequential KDSL (2025a) Qwen2-VL 2B Qwen2-VL 7B SLM rules; LLM inference
Sequential TrajAgent (2024) TrajFormer GPT-4 TaxiBJ, Porto, etc. LLM plans; SLMs execute
II. Cost-effective Collaboration
Cascade R2R (2025a) DeepSeek-R1-1.5B DeepSeek-R1-32B AIME, GPQA Token-level routing for efficiency
TTS-optimal AgentTTS (2025b) LLaMA-3-3B LLaMA-3-70B HotpotQA, CWQ, etc. Budget-aware collaboration
III. Privacy-aware Cloud–Edge Collaboration
SLM learner MiniLLM (2024) GPT-2-125M GPT-2-1.5B DollyEval, SelfInst, etc. Distill LLM into SLM
LLM learner Proxy-tuning (2024a) LLaMA-2-7B LLaMA-2-70B GSM, Toxigen, etc. Steer LLM with private SLM
IV. Trustworthiness Collaboration
Safe decoding LLM Purifying (2024) CodeLLaMA-7B CodeLLaMA-13B HumanEval, LAMBADA, etc. Logit fusion for clean gen.
SLM guard STAND-Guard (2024) Mistral-7B GPT-3.5/4 BeaverTails, SafeRLHF Task-adaptive moderation

### v2

@@ 4.2 Fine-tuning Stage
During fine-tuning, edge data can facilitate the evolution of both edge and cloud LLMs by providing valuable fine-tuning signals. However, ensuring data protection without exposing sensitive information to the cloud remains a major challenge. To address this, privacy-preserving edge–cloud training frameworks have been proposed, which can be categorized into three paradigms based on the learning target: (1) SLMs as Learners: the edge SLM acquires new capabilities from larger LLMs through local fine-tuning, distillation, or personalization, guided by cloud LLMs without revealing raw data (Fig. 15); (2) LLMs as Learners: the cloud LLM improves by consuming proxy signals or adapters distilled from locally trained SLMs, thereby enhancing generalization without accessing private data (Fig. 16); and (3) Collaborative Learners: both SLMs and LLMs co-evolve through federated or privacy-preserving coordination (Fig. 17).
@@
(i) Raw Data: SLMs enhance reasoning by learning from cloud LLM outputs while keeping private data local. The LLM provides interpretable supervision (e.g., responses, reasoning traces) or synthetic samples for SLM fine-tuning. MiniLLM Gu et al. (2024) and LlamaDuo Park et al. (2025) distill reasoning traces, while DRAG Chen et al. (2025c) extends to retrieval-augmented reasoning. Privacy-aware frameworks such as HomeLLaMA Huang et al. (2025c) and  Qin et al. (2024) use selective textual supervision to align SLMs with user preferences under strict privacy constraints.
(ii) Logit-based Learning: SLMs learn from teacher LLMs’ numerical outputs, using logits as soft targets to imitate their decision boundaries without accessing raw text. TinyLLM Tian et al. (2025) distills from multiple teachers for robustness, Mix Distillation Li et al. (2025g) combines LLM and SLM supervision, and ADPA Gao et al. (2025b) introduces an advantage-based objective for better reasoning alignment, ensuring privacy and efficiency through abstracted signal exchange.
@@
(i) Logit-based Learning: Instead of directly fine-tuning an LLM, Ormazabal et al. (2023) demonstrate that an LLM can be steered by tuning an SLM on private data and exposing only the SLM’s output logits to guide the LLM’s responses. Specifically, they train a lightweight MLP on the logits produced by both the SLM and the LLM, which is then used to steer the LLM’s output during inference. This approach enables the use of private data for adaptation while avoiding its direct exposure to the LLM. An extension of this method involves applying direct logit offsets between the tuned model and the LLM Liu et al. (2024a); Wang et al. (2025d), by computing the logits difference between the trained SLM and its frozen counterpart. The LLM adds this offset to its own output logits during inference.
Follow-up work has examined the limitations of this approach. CPT He et al. (2024) and GradOT Yao et al. (2025) address inconsistencies arising from discrepancies between training and inference pipelines. Whereas the SLM is trained alone, it is expected to interact with its frozen counterpart to compute the logits offset and steer the LLM during inference. To address this gap and improve alignment, CPT integrates both the frozen SLM and LLM in the training process to closely mirror the inference phase interaction and include the LLM feedback in the gradient updates. On the other hand, GradOT approximates the gradient updates using offsite-tuning Xiao et al. (2023) to enable efficient alignment without co-training. To enhance robustness, GOOD Fang et al. (2024a) decodes the SLM’s output into text and re-encodes it for the LLM to extend logit steering to SLM–LLM pairs without a shared vocabulary.
(ii) LoRA-based Learning: Recent work leverages LoRA signals to propagate knowledge instead of requiring logit outputs from SLMs to enable adaptation without direct access to private data. LoRASuite Li et al. (2025f) and LoRA-X Farhadzadeh et al. (2025) demonstrate how LoRA signal transfer can be applied across model versions and architectures. Similarly, Trans-LoRA Wang et al. (2024c) and Cross-LoRA Xia et al. (2025) extend this capability to heterogeneous model families, enabling more generalizable and privacy-conscious adaptation of LLMs.
@@
Beyond one-way transfer, collaboration can be bidirectional: SLMs and LLMs co-adapt via shared representations, logits, or adapter weights while preserving privacy. Fig. 17 shows multiple edge SLMs learn from local private data and transmit abstracted updates, such as logits or LoRA adapters, to the cloud LLM for aggregation and refinement, and then the cloud LLM returns high-level knowledge or refinement feedback to the local. CrossLM Deng et al. (2025) achieves mutual enhancement by combining private task data and LLM-generated samples for privacy-preserving co-training. LSRP Zhang et al. (2025c) enables bidirectional feedback, where the LLM guides SLM retrieval and refines its own responses. Federated frameworks such as FedMKT Fan et al. (2025b), FedPT Gao et al. (2024), FedPFT Peng et al. (2024), FDLoRA QI et al. (2024), and the Automated Federated PEFT Pipeline Fang et al. (2024b) aggregate client-side SLM adapters or logits to update central LLMs while preserving data privacy. FLoRA Wang et al. (2024e) further improves personalization through enhanced edge aggregation.
@@ 5 Collaboration for Trustworthiness
LLMs offer strong generative ability and broad knowledge coverage but face major trustworthiness challenges in real-world use. Uncontrolled training data can lead to hallucinations, harmful or biased outputs, privacy leaks, and vulnerabilities exploitable through jailbreaks or adversarial attacks Zhang et al. (2025d); Liu et al. (2023); Wu et al. (2025); Zhang et al. (2025b); Liu et al. (2020); Wang et al. (2021). Updating LLMs for safety is costly and may harm generation quality. A more practical solution is integrating safety policies that guide generation without retraining. SLMs serve this role effectively due to their compactness, independence from LLM architectures, and ease of policy updates. Recent SLM–LLM collaborations use SLMs to encode or enforce safety policies while LLMs handle content generation. Two paradigms dominate: (1) Safety-guided decoding, where an SLM modifies LLM logits during decoding for token-level control; and (2) Guardian–generator, where the SLMs trained as Guardians filter inputs and audit outputs, providing interaction-level governance. These paradigms are shown in Figs. 18 and 19.
@@ 5.1 Safety-Guided Decoding
Unsafe outputs in LLMs often arise during decoding; thus, integrating safety policies into token-level generation is crucial yet challenging, as it is difficult to anticipate which tokens may trigger unreliable content. Recent studies employ safety-tuned SLMs, fine-tuned for risks such as jailbreaks, privacy leaks, or bias, to guide LLM decoding in a safety-aware manner. As shown in Fig. 18, SLM and LLM logits are fused to ensure both quality and safety through two strategies: (1) direct logit fusion (linear combination) and (2) logit offset fusion (using SLM logits as offsets to steer LLM probabilities toward safer tokens).
@@
(1) Direct Logit Fusion
The CP-$\Delta$ theory Vyas et al. (2023) provides an information-theoretic foundation for model fusion, demonstrating that combining the predictive distributions of two models, $q_{1}(y|x)$ and $q_{2}(y|x)$, with distinct copyright risks yields a fused distribution $p(y|x)\propto q_{1}(y|x)^{\alpha}q_{2}(y|x)^{1-\alpha}$ that minimizes weighted KL divergence and balances their strengths without retraining, where $\alpha$ controls each model’s contribution. The fused distribution lies closer to both sources than they are to each other, thus achieving a purification effect, retaining desirable properties while reducing copyrighted content. This operation is equivalent to linear weighting in logit space, $z_{p}=\alpha z_{1}+(1-\alpha)z_{2}$, providing the theoretical foundation for SLM–LLM collaborative logit fusion. Subsequent methods build on this principle: Purifying LLMs Li et al. (2024a) fuses clean SLM and contaminated LLM logits to mitigate copyright, privacy, and backdoor risks, while MOD Shi et al. (2024) extends fusion to multiple objectives, integrating models optimized for fluency, faithfulness, and safety to enable controllable generation.
(2) Logit Offset Fusion
Mitchell et al. (2024) introduced the concept of logit offset fusion by decomposing a fine-tuned model’s output into the base model’s logits and a behavioral delta:
$z_{\text{ft}}(x)=z_{\text{base}}(x)+\Delta(x),$ (1)
where $\Delta(x)$ represents the behavioral delta induced by fine-tuning. Through an up-scaling strategy, it integrates a large pre-trained model’s logits $z_{\text{large}}$ with the behavioral delta of a small fine-tuned model $\Delta(x)=z_{\text{ft}}(x)-z_{\text{small}}(x)$:
$z_{\text{fusion}}(x)=z_{\text{large}}(x)+\lambda\,\Delta(x),$ (2)
where $\lambda$ is a scaling factor aligning the magnitude of the small model’s delta with the large model’s logit scale. This simple logit-level combination achieves up to 73% of the full large-model fine-tuning helpfulness without re-training the large model. This demonstrates that safety-aligned small models’ behavioral deltas can enhance LLM safety when fused at the logit level. Building on this idea, Offset Unlearning Huang et al. (2025a) trains SLMs to generate targeted logit offsets that suppress sensitive or fabricated knowledge for privacy-preserving unlearning. Conversely, Weak-to-Strong Jailbreak Zhao et al. (2025) reveals that unsafe SLMs can steer safe LLMs toward harmful outputs, implying that the same mechanism can defensively employ safe SLMs to counter jailbreak attacks through safety-guided logit fusion.
@@ 5.2 SLM as Guardian, LLM as Generator
Figure 19: SLM Guardian, LLM Generator.
Building on this paradigm, later studies advanced the framework. Llama Guard Inan et al. (2023) built a bidirectional safety pipeline using a fine-tuned LLaMA-2-7B classifier. Llama Guard 2 Team (2024) added multi-label risk classification; Llama Guard 3 Dubey et al. (2024) expanded multilingual support and inspired lightweight variants like Llama Guard 3-1B-INT4 Fedorov et al. (2024) and cross-modal Llama Guard 3 Vision Chi et al. (2024). Llama Guard 4 AI (2025a) further refined the taxonomy to 14 safety categories and was adapted for healthcare auditing Gangavarapu (2024). For lightweight protection, Prompt Guard AI (2024) and Prompt Guard 2 AI (2025b), based on DeBERTa-v2 He et al. (2021), serve as fast, low-cost input filters. ShieldGemma Zeng et al. (2024, 2025) and WildGuard Han et al. (2024) use synthesized, large safety data to boost robustness against adversarial prompts. ThinkGuard Wen et al. (2025) adds “slow thinking” for better interpretability and defense. Beyond text safety, MiniCheck Tang et al. (2024) applies the guardian concept to factual verification, while Cascade Guardrails Nagireddy et al. (2024) explores hierarchical defenses. LlamaFirewall Chennabasappa et al. (2025) extends this paradigm to multi-agent settings.
@@ 5.3 Discussion
6 Applications
Collaboration for performance
In guidance-generation, SLMs assist LLMs in reasoning and classification tasks, such as intent detection Liang et al. (2024), natural language inference Xu et al. (2024a), sentiment analysis Mojarradi et al. (2024), and multi-hop question answering Lee et al. (2024). In division–fusion, they form multi-stage pipelines for domains such as clinical analysis Gondara et al. (2025), factual verification Hsu et al. (2024), NL2SQL Fan et al. (2024), and trajectory modeling Du et al. (2024), where SLMs handle subtasks and LLMs integrate the results. This suggests that guidance-enhanced generation is more suitable for common general-domain tasks, where LLMs are the backbone generator and SLMs are trained on local downstream data, while division–fusion collaboration is preferred when both models exhibit comparable capabilities across different subtasks.
Collaboration for cost-effectiveness
The cost-effective SLM–LLM collaboration methods show stage-specific application preferences. Pre-training-stage frameworks target corpus construction and scaling efficiency on large datasets (e.g., Pile, FineWeb-Ed, C-Eval), supporting sustainable and scalable model development Rawat et al. (2024); Yano et al. (2025); Goyal et al. (2025). Fine-tuning-stage methods focus on intent classification and reasoning transfer (e.g., BANKING, CLINC, MMLU, GSM8K), enabling smaller models to inherit LLM capabilities with minimal cost Jazbec et al. (2024); Zhang et al. (2025a); Fan et al. (2025c). Inference-stage approaches emphasize real-time reasoning and conversational tasks (e.g., AIME, GPQA, ShareGPT) where latency and compute efficiency are critical Chen et al. (2024a); Ong et al. (2025); Zhang et al. (2024d).
Collaboration for cloud-edge privacy
The inference-stage methods are particularly suited for end-user scenarios where privacy is a main concern. Edge SLMs prevent leakage of PII or sensitive context to external providers Chong et al. (2024); Zhang et al. (2024f), supporting domains such as healthcare and customer service. Depending on privacy needs, SLMs as Information Gatekeepers may be applied Zhang et al. (2024i); Cheng et al. (2025). Meanwhile, proxy tuning enables domain-specific or personalized LLM adaptation without exposing private data, offering broader potential under edge resource privacy constraints and edge SLMs’ limited reasoning and context capacity Li et al. (2025f); Wang et al. (2025d); Deng et al. (2025).
Collaboration for trustworthiness
SLM–LLM collaboration for trustworthiness is applied in safety-critical and compliance-sensitive contexts. Safe decoding methods can correct hallucination, privacy, and toxicity risks in code generation, instruction following, and dialogue by steering the decoding distributions Li et al. (2024a, a); Mitchell et al. (2024); Huang et al. (2025a). Guardian frameworks support content moderation, assistants, and fact verification by filtering directly unsafe inputs and ensuring output reliability Inan et al. (2023); AI (2024); Zeng et al. (2024); Wen et al. (2025). This suggests that the former is better suited for applications requiring a balance between output quality and truthfulness, whereas the latter is preferred for highly secure applications that eliminate potentially risky inputs and outputs.
@@ 7 Challenges and Future Directions
Cost-Aware SLM–LLM Collaboration Current SLM–LLM collaborations emphasize performance over cost modeling, risking inefficient resource use. Future work should develop adaptive orchestration frameworks to optimize cost–performance trade-offs through predictive cost-aware routing, non-myopic collaboration policies considering long-term budgets, and multi-objective optimization balancing accuracy, latency, and financial cost for sustainable, scalable SLM–LLM systems.
Theory for Privacy-Preserving Collaboration Current privacy-preserving SLM–LLM collaborations lack formal theoretical foundations, relying mostly on empirical evaluations without quantifiable privacy assurances. Future research should establish rigorous metrics and provable bounds for data leakage and privacy preservation, enabling verifiable guarantees of data security and fostering trust in privacy-sensitive SLM–LLM systems.